Help me with the workings to these questions please!!

Question 8 Not answered Marked out of 1.00 A physical-fitness center has found that, from those members who perform high-impact exercising on one visit, 96% will do the same on the next visit, while on those who perform low-impact exercising on one visit, 78% will do the same on the next visit. On the last visit, suppose 30% of members did high-impact exercising. After two more visits, the percentage of members performing low-impact exercising will be: Flag question Give you answer correct to two decimal places.) Answer The correct answer is: 45.29 Question 9 Not answered Marked out of 1.00 In Australia, 63% of the population are rugby fans. Among those who are rugby fans, 64% are also soccer fans. Among those who are not rugby fans, 89% are soccer fans. Julie is a soccer fan, and her probability of being a rugby fan is Give you answer correct to three decimal places.) Flag question Answer: The correct answer is: 0.55 Question 10 Not answered Marked out of 1.00 A financial advisor wants to create a portfolio consisting of stocks and bonds, with exactly 7 bonds and double the number of stocks. If s-27+4 stocks and b-7+2 bonds are acceptable for the portfolio, the number of ways in which the portfolio could be created is: Flag question Give you answer correct to three decimal places.) Answer: The correct answer is: 110160.00
Answer #1

Q8) If 30% did high impact exercising , then the percentage of people from this 30% who will do high impact exercising next day= 0.96*30% = 28.8%

And 70% did low impact exercising , then the percentage of people from this 70% who will do high impact exercising next day= 0.22*70% = 15.4%

So, total population doing high impact exercising is = 28.8+15.4= 44.2%

Out of these 44.2% population who did high impact exercising, the percentage of people who will do high impact exercising the 2nd day will be = 0.96*44.2 = 42.43%

And out of the remaining 55.8% population who did low impact exercising, the percentage of people who will do low impact exercising the 2nd day = 0.22*55.8% = 12.28%

So total percentage of people doing high impact exercising after 2nd day = 42.43+12.28 = 54.71%

So total percentage of people doing low impact exercising after 2nd day = (100-54.71) = 45.29%

Q9) Percentage of people who are rugby fans as well as soccer fans= 63%*0.64 = 40.32%

Percentage of people who are not rugby fans but are soccer fans= 37%*0.89 = 32.9%

So total percentage of population who are soccer fans = 40.32+32.9 = 73.22%

So percentage of soccer fans who are rugby fans= 40.32/73.22 = 55.06%

So the probability of Julie who is a soccer fan , is also a rugby fan = 0.55.

Q10) Number of bonds to be selected = 7

Number of stocks to be selected = 7*2 = 14

Total number of stocks availabe for selection = 2*7+4 = 18

Total number of bonds available for selection = 2+7 = 9

So, number of ways in which portfolio can be created = (18C14)*(9C7) = 110160.000